# Efficient Reasoning for Large Reasoning Language Models via Certainty-Guided Reflection Suppression

### v1

@@ 4.2 Comparison with State-of-the-art Methods
In Table 1, we compare the proposed CGRS method with previous efficient reasoning methods in terms of accuracy and output length. The evaluation covers diverse model scales and architectures, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, QwQ-32B, and Qwen3-4/8/14/32B across four reasoning benchmark datasets (AIME24, AMC23, MATH500, and GPQA-D). The results for Qwen3-32B are provided separately in the Appendix.
As can be seen, CGRS consistently achieves the optimal balance between average length reduction ratio and average accuracy preservation across diverse models, demonstrating its effectiveness for efficient reasoning. For example, on the Qwen3-14B model, CGRS reduces output length by $25.4\%$ while maintaining accuracy. In contrast, while baseline methods achieve greater length reduction ($77.3\%$ for NoThinking and $38.8\%$ for DEER), they suffer significant accuracy degradation ($17.8\%$ and $5.7\%$ drops, respectively). TALE maintains comparable accuracy but achieves less length reduction than CGRS ($16.2\%$ vs. $25.4\%$).
Moreover, the proposed CGRS method demonstrates reliable effectiveness in every test case, consistently achieving $18.5\%$ to $41.9\%$ length reduction with negligible accuracy drop (up to $2\%$), while the baseline methods often perform unsatisfactorily. For example, NoThinking prompts models to bypass slow thinking processes, which yields high compression rates but severely degrades accuracy (e.g., $65.4\%$ length reduction and $9.4\%$ accuracy drop on MATH500 with DeepSeek-R1-Distill-Qwen-7B). Early-exit approaches like Dynasor and DEER also sometimes underperform, e.g., on AMC23 with DeepSeek-R1-Distill-Llama-8B, they only achieve $10.9\%$ and $14.0\%$ compression, respectively, with DEER additionally reducing accuracy. Remarkably, in the same test case, CGRS attains $23.1\%$ length reduction while slightly improving accuracy.
Notably, on the QwQ-32B model, the baseline methods (TALE, NoThinking, Dynasor, and DEER) achieve only minor reduction ratios across all four benchmark datasets. This limitation occurs because these methods rely on hard-coded reflection boundaries that assume reasoning terminates at the </think> token, while QwQ-32B often continues reflecting beyond this token (Yang et al. 2025b). In contrast, the proposed CGRS method dynamically regulates reflection triggers based on confidence scores, which is independent on the end-of-think token, enabling CGRS to maintain strong compression rates ($30.5\%$) while preserving accuracy in this case.

### v2

@@ 4.2 Comparison with State-of-the-art Methods
In Table 1, we compare the proposed CGRS method with previous efficient reasoning methods in terms of accuracy and output length. The evaluation covers diverse model scales and architectures, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, QwQ-32B, and Qwen3-4/8/14/32B across four reasoning benchmark datasets (i.e., AIME24, AMC23, MATH500, and GPQA-D).
As can be seen, CGRS consistently achieves the optimal balance between average length reduction ratio and average accuracy preservation across diverse models, demonstrating its effectiveness for efficient reasoning. For example, on the Qwen3-14B model, CGRS reduces output length by $25.4\%$ while maintaining accuracy. In contrast, while baseline methods achieve greater length reduction ($77.3\%$ of NoThinking and $38.8\%$ of DEER), they suffer significant accuracy degradation ($17.8\%$ and $5.7\%$ drops respectively). TALE maintains comparable accuracy but achieves less length reduction than CGRS ($16.2\%$ vs. $25.4\%$).
Moreover, the proposed CGRS method demonstrates reliable effectiveness in every test case, consistently achieving $18.5\%$ to $41.9\%$ length reduction with negligible accuracy drop (up to $3\%$), while the baseline methods often perform unsatisfactorily. For example, NoThinking prompts models to bypass slow thinking processes, which always yields high compression rates but severely degrades accuracy (e.g., $65.4\%$ length reduction and $9.4\%$ accuracy drop on MATH500 with DeepSeek-R1-Distill-Qwen-7B). Early-exit approaches like Dynasor and DEER also sometimes underperform, e.g., on AMC23 with DeepSeek-R1-Distill-Llama-8B, they only achieve $10.9\%$ and $14.0\%$ compression, respectively, with DEER additionally reducing accuracy. Remarkably, in the same test case, CGRS attains $23.1\%$ length reduction while slightly improving accuracy.
Notably, on the QwQ-32B model, the baseline methods (TALE, NoThinking, Dynasor, and DEER) achieve only minor reduction ratios across all four benchmark datasets. This limitation occurs because these methods rely on hard-coded reflection boundaries that assume reasoning terminates at the </think> token, while QwQ-32B often continues reflection behaviors beyond this token (yang2025dynamicearlyexitreasoning). In contrast, the proposed CGRS method dynamically regulates reflection triggers based on confidence scores, which is independent on the end-of-think token, enabling CGRS to maintain strong compression rates ($30.5\%$) while preserving accuracy in this case.
@@ 5 Conslusion
Acknowledgments
This work is supported in part by the State Key Laboratory of General Artificial Intelligence. DH is supported by National Science Foundation of China (NSFC62376007) and Beijing Natural Science Foundation (Z250001).
Appendix A Additional Details of Reflection Triggers Selection in Section 3.3
As described in Section 3.3, we focus on four primary trigger words (“Wait”, “But”, “Alternatively”, and “Hmm”) in our experiments. For each trigger word, we identify all natural language variants present in the tokenizer’s vocabulary. For example, the keyword “Wait” has variants such as “.Wait,” “␣WaitForSeconds,” “.wait,” “.WaitFor,” “ĉwait,”. Then we analyze their generation frequencies in reasoning, in order to filter out tokens that are unlikely to appear (e.g., “.Wait”) during actual reasoning. This analysis uses reasoning traces from the DeepSeek-R1-Distill-Qwen7B model on both the AIME24 and AMC23 benchmarks. We examine two tokenizer types: LlamaTokenizerFast (for DeepSeek-R1-Distill-Llama-8B) and Qwen2Tokenizer (for DeepSeek-R1-Distill-Qwen-7B, the Qwen-3 family, and QwQ-32B). The complete list of filtered trigger words (including variants) and their corresponding token IDs are provided in Table 3.
Appendix B Examples of Redundant Reflection Behaviors
In this section, we present qualitative examples demonstrating redundant reflection behaviors in LRLM reasoning. We analyze responses from the DeepSeek-R1-Distill-Qwen-7B model on a complex combinatorial problem in the AIME24 dataset.
As shown in Table 5, the model initially identifies a correct or promising approach but subsequently enters a prolonged reflective phase. This phase is characterized by ineffective self-correction, repeated attempts at previously explored approaches, and overthinking trivial details—often initiated by reflection triggers (like “Wait”, “But” and “Alternatively”). Despite the extended reasoning, these additional steps fail to improve answer quality and, in some cases, introduce new errors or overwrite correct logic. These failure cases underscore the need for dynamic reflection control during inference. By identifying the onset of these inefficient reflective loops, CGRS helps maintain concise and effective reasoning without suppressing productive analysis.
Trigger Words Qwen Token ID Llama Token ID
Wait 14190 14524
␣Wait 13824 14144
wait 11489 11748
␣wait 3783 3868
But 3983 4071
␣But 1988 2030
but 8088 8248
␣but 714 719
Alternatively 38478 93114
␣Alternatively 38478 39578
Alternative 75763 76863
␣Alternative 41109 42209
Hmm 80022 81122
␣Hmm 88190 89290
Table 3: Reflection trigger words and their corresponding token IDs in Qwen2Tokenizer and LlamaTokenizerFast tokenizers. ␣ denotes a space character.
Appendix C Ablation on Threshold $\delta$
The results are shown in Table 4. As $\delta$ decreases, the model becomes less strict in suppressing uncertain reasoning steps, leading to longer responses but lower accuracy. Specifically, when $\delta$ drops from 0.9 to 0.1, accuracy decreases from 88.3% to 72.5%, while the length reduction (LR) weakens from 41.9% to 53.7%. This demonstrates the trade-off controlled by $\delta$: higher thresholds enforce more confident reasoning and concise outputs, whereas lower thresholds allow more exploratory but error-prone reasoning steps.
[…]
0.9 88.3% 41.9
0.5 76.2% 48.5
0.1 72.5% 53.7
Table 4: Effect of threshold $\delta$ on accuracy and length reduction (LR) on AMC23 and R1-Distill-Qwen-7B.
Reflection Examples Analysis
Case 1: Failed self-correction. “Therefore, $r(S)$ is $S^{\prime}$ which is equivalent to $S\cap r^{-1}(S^{\prime})=S$. [Correct] Wait, maybe that’s not the right way. Alternatively, since $r$ is a permutation, $r(S)$ is just another set of the same size as $S$. So, if $S^{\prime}$ is a subset of $S$, then $r(S)$ must be a subset of $S^{\prime}$ which would require that $S^{\prime}$ is a subset of $r^{-1}(S^{\prime})$. But since $r$ is a rotation, it’s a bijection, so $r^{-1}(S^{\prime})$ is just another set of size $|S|$. [Wrong] correct to wrong: The initial thought process was on the right track, but the self-correction led to incorrect reasoning steps.
Wait, but perhaps the number is 105. [Wrong] Wait, but 105 is $2\cdot 10^{2}$, which is 105. Wait, perhaps the number is 105. [Wrong] wrong to wrong: Repeated incorrect self-correction, circling back to the same incorrect value.
Case 2: Reattempt a previously tried approach. Wait, but in this case, it’s more specific: the set must be entirely disjoint from some rotated copy of itself. Hmm. Wait, maybe it’s useful to think in terms of necklace structures. A necklace is an equivalence class under rotation, but I’m not sure. [The method “necklace counting” was tried before and failed.] This highlights an instance where the model redundantly reapplied a previously attempted strategy it had already recognized as ineffective for the current problem.
Case 3: Overthinking trivial details. Wait, but actually, for the identity rotation, $r(S)=S$, so $S\cap r(S)=S$, which is not empty unless $S$ is empty. So, for the identity rotation, the condition $S\cap r(S)=\emptyset$ is only satisfied when $S$ is empty. So, for non-empty $S$, the identity rotation cannot satisfy the condition. Therefore, for non-empty $S$, we need some non-identity rotation $r$ where $S\cap r(S)=\emptyset$. Wait, actually, the problem says ‘can be rotated’, so it includes the identity rotation. But as we saw, for non-empty $S$, the identity rotation only works if $S$ is empty, which is trivial. So, perhaps the problem is really about non-empty $S$ where some non-identity rotation makes $S\cap r(S)=\emptyset$.” This lengthy deliberation focuses on the trivial case of the identity rotation and its implications. This demonstrates a scenario where the model’s reflection-driven over-complication of a simple problem result in unnecessarily verbose reasoning.
Table 5: Examples of redundant reflection behaviors extracted from the response generated by the DeepSeek-R1-Distill-Qwen-7B model on a specific problem in AIME24 dataset. The original problem is “Each vertex of a regular octagon is independently colored either red or blue with equal probability. The probability that the octagon can then be rotated so that all of the blue vertices end up at positions where there were originally red vertices is $\tfrac{m}{n}$, where $m$ and $n$ are relatively prime positive integers. What is $m+n$?
Appendix D Relationship Between Certainty and Correctness
Figure 4: Certainty distributions of probes with correct and incorrect results on AIME24, AMC23, and MATH500.
Our further analysis correlation between reasoning certainty and correctness. Specifically, we collect reasoning trajectories generated by DeepSeek-R1-Distill-Qwen-7B on three benchmarks: AIME24, AMC23, and MATH500. For each trajectory, we insert fixed-interval probing points every 128 tokens to examine (1) the correctness of the probed content, and (2) the certainty score computed by our method. As shown in Figure 4, reasoning steps with high certainty (e.g., >0.9>0.9) are significantly more likely to be correct, indicating that the model’s self-estimated certainty serves as a reliable indicator of reasoning quality.